Sales                            $300,000

Variable Costs             $180,000

Fixed Expenses            $70,000

Units sold                    20,000 units

What is the breakeven in units?

How many units must be sold to reach a target profit of $100,000 before taxes?

If sales increased by 5%, by what % would profits increase?
Answer #1

Selling price per unit = Sales/Units

= 300,000/20,000

= $15

Variable cost per unit = Variable cost/units

= 180,000/20,000

= $9

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 15-9

= $6

Break even point ( in units) = Fixed Expenses/Contribution margin per unit

= 70,000/6

= 11,667 units

Number of units to be sold too earn target profit = ( Fixed Expense + Target profit) / Contribution margin per unit

= ( 70,000+100,000)/6

= 28,334 units

Sales 300,000
Variable costs -180,000
Contribution Margin 120,000
Fixed expenses -70,000
Operating Profit $50,000

Operating leverage = Contribution margin/Operating income

= 120,000/50,000

= 2.4

Increase in profit = Operating leverage x Increase in sales

= 2.4 x 5%

= 12%

Kindly comment if you need further assistance. Thanks‼!